OpenStudy (anonymous):

which of the following are square roots of the number below ? 81 a. 81^1/2 b.18 c.9 d.-(81) ^1/2 ?

OpenStudy (apoorvk):

sqrt of a no. is that, which when when multiplied by itself, gives the original no. so if 'a' is a sqrt of 'b', the a^2=b now what do you think would the answer be??

OpenStudy (anonymous):

c.

OpenStudy (anonymous):

a too

OpenStudy (apoorvk):

and, D too. can you think why? so its ACD

OpenStudy (anonymous):

d is not

OpenStudy (anonymous):

\[\sqrt{-81}= i \sqrt{81}\] square this and u get -81

OpenStudy (apoorvk):